OpenStudy (anonymous):

What is the sum of the arithmetic sequence 174, 168, 162 …, if there are 33 terms? 2,541 2,574 2,607 2,640

OpenStudy (anonymous):

a=174 c.d. d=168-174=-6 n=33 \[Sn=\frac{ n }{ 2 }\left\{ 2a+\left( n-1 \right)d \right\}\] find S33.

OpenStudy (anonymous):

Please help with a little more im getting a decimal

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[S33=\frac{ 33 }{2 }\left\{ 2*174+\left( 33-1 \right)\left( -6 \right) \right\}\] \[=33\left\{ 174+32\left( -3 \right) \right\}=33\left( 174-96 \right)=33*88=?\]

OpenStudy (anonymous):

2,904

OpenStudy (anonymous):

correction 174-96=78 not 88 33*78=?

OpenStudy (anonymous):

2,574

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw